# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 4.3. Numerics
Figure 4. Trajectories $x(\cdot;x^{0},u)$ of system Equation 30 and its kEDMD surrogate $\widehat{x}(\cdot;x^{0},u)$ generated with Algorithm 1 (left) and their deviation $\|x(k;x^{0},u)-\widehat{x}(k;x^{0},u)\|_{2}$, $k\in[1:50]$, in norm (middle) for initial value $x^{0}=(0.1,0.1)^{\top}$ for the control sequence $u$ (right).
Next, we randomly chose 20 initial values $\hat{x}_{i}$ in $[-0.5,0.5]^{2}\subset\Omega$, as well as 20 random input sequences $\{u_{i}\}\subset[-0.1,0.1]\subset\Omega$ of length 30. These smaller boxes ensure that the trajectories remain in the domain $\Omega$.
In Figure 5 we depict statistical information on the error. The areas between maximum and median of the errors for the different numbers of cluster points $d$ barely overlap, that is, the reduction of the errors in the number of macro points $d$ is well noticeable. For $d=441$ the maximal error exceeds the $10^{-1}$ threshold nearly immediately after a few time steps and even the median transcends this bound at about 15 time steps are conducted. The greater $d$ is selected, the smaller the errors turn out.
Figure 5. Median and maximum of error of the kEDMD-based surrogate model Equation 25 to the real model Equation 30 from 20 trajectories for $d=1681$ cluster points. In between the median and maximum of the errors, the $0.8,0.9,0.95$-quantiles are plotted (left). Median and maximum for $d=441,1681,6561$ cluster points in one plot (right).
Finally, in Figure 6, we inspect the error of a one-step prediction in the Euclidean norm, similar to Figure 1 in Section 3.3. The learning of the kEDMD-based system is again performed for the three numbers of cluster points $d=441,1681,6561$ for the Chebyshev-based macro grid and $N=25$ data points in a neighborhood of $\varepsilon=\frac{1}{d}$. The errors $\max_{j\in[1:20]}\|F(x,u_{j})-\widehat{F}(x,u_{j})\|$ on a respective validation grid in $[-1,1]^{2}$ are computed for the control values in $\{-2+0.2(j-1)|\ j\in[1:20]\}$. In the intensity plots of Figure 6, it can be observed that the error evenly decreases the greater $d$ is chosen. This validates our results from Theorem 4.3: There are two sources of error that can be influenced, i.e., the fill distance of the macro grid $h_{\mathcal{X}}$ depending on number of cluster points $d$ and the maximal cluster size $\varepsilon$. If $\varepsilon$ is small enough to compensate the factor $\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}$ that increases with $d$, the decrease of the fill distance by refining $\mathcal{X}$ yields a decay of the approximation error.
Figure 6. Approximation error $\max_{j\in[1:20]}\|F(x,u_{j})-\widehat{F}(x,u_{j})\|$ for the predicted system Equation 30 with $d\in\{441,1681,6561\}$ cluster points using a Chebychev-based grid (from left to right).
@@ 5. Conclusions and outlook
In this work, we provided a novel kernel EDMD scheme with flexible state-control sampling and uniform error estimates for data-driven modeling of dynamical (control) systems with stability guarantees.
In the first part of this work, we extended existing uniform bounds [20] on the full approximation error on Koopman approximants generated with kernel EDMD to Tychonov-regularized ones and derived proportional error bounds. While the first is important for robustness in view of the poor conditioning of kernel matrices (e.g. for noisy data), the second is key to show that asymptotic stability of an equilibrium w.r.t. the surrogate dynamics, certified by a Lyapunov function, is preserved for the original dynamical system and vice versa.
In the second part, we proposed a novel kernel-EDMD scheme for control-affine systems building upon arbitrarily-sampled control-state pairs and rigorously showed the first uniform (proportional) bounds on the full approximation error for control systems. We demonstrated that, as a consequence, stabilizing feedback laws designed for the data-driven surrogate models also ensure asymptotic stability w.r.t. the closed-loop of the original system building upon the proposed stability-analysis framework for kernel EDMD. We accompanied our findings by various numerical simulations.
Future work will be devoted to data-driven controller design leveraging the proposed highly-flexible sampling regime in combination with the novel proportional bounds derived in a non-restrictive setting. Herein, results from [12] and [6] may be leveraged for robustification and direct data-driven controller design.
@@ A.1. Error bounds
The first claim on the error between $\mathcal{K}$ and $\widetilde{\mathcal{K}}_{\lambda}$ follows by a combination of [20, Theorem 5.2] and Lemma 2.5. The proof of the second claim on $F-\widetilde{F}_{\lambda}$ is very similar to that of Corollary 2.8. ∎
@@ Appendix B The regularization operator
[…]

### v3

@@ 4.3. Feedback stabilization
Figure 5. Left: KEDMD approximants $\widehat{x}(\cdot;x^{0},0)$ (open loop) and $\widehat{x}_{\kappa}(\cdot;x^{0})$ (closed loop) of system eq. 31, initial value $x^{0}=(0.5,0.1)^{\top}$, generated with algorithm 1 based on the Padua grid with $d=435$ cluster points. Deviations $\|x(k;x^{0},0)-\widehat{x}(k;x^{0},0)\|_{2}$ and $\|x_{\kappa}(k;x^{0})-\widehat{x}_{\kappa}(k;x^{0})\|_{2}$, $k\in[1:100]$, in open-(mid) and closed-loop (right) control for $d\in\{435,1653,6555,11476\}$.
We first inspect the open-loop predictions for the uncontrolled case. As can be seen in Figure 5 (left), starting at the initial state $x^{0}=(0.5,0.1)^{\top}$ the behavior of the ground truth is accurately replicated, as both trajectories move towards the second equilibrium at $x^{*}=(0.55,0.66)^{\top}$ illustrating that this initial value is not contained in a potential domain of attraction. The kEDMD-based surrogate can only approach $x^{*}$ up to a certain accuracy, as shown in the error plots (Figure 5, mid). It can be observed, that the error decreases the greater the number $d$ of Padua grid points is. Furthermore, Figure 5 exhibits the results for using the nominally stabilizing feedback $\kappa$. The feedback law also stabilizes the origin w.r.t. the kEDMD approximants as can be inferred from the numerically computed trajectories of system eq. 31, which tend towards the origin. When using $\kappa$ also for the kEDMD-based surrogate systems, we observe the statement of Proposition 4.6 (with switched roles of surrogate and ground truth). The trajectories approach a neighborhood of the origin depending on the number of cluster points. These results confirm that asymptotically stabilizing feedbacks for the original system are practically asymptotically stabilizing w.r.t. the kEDMD-based surrogate. Future work considers leveraging the converse statement (as proven in Proposition 4.6) for data-driven computation of feedback laws.
@@ 5. Conclusions and outlook
We provided a novel kernel EDMD scheme with flexible state-control sampling and uniform error estimates for data-driven modeling of dynamical (control) systems with stability guarantees. In the first part of this work, we extended existing uniform bounds [28] on the full approximation error on Koopman approximants generated with kernel EDMD to regularized ones and derived proportional error bounds. While the first is important for robustness in view of the poor conditioning of kernel matrices (e.g. for noisy data), the second is key to show that asymptotic stability of an equilibrium w.r.t. the surrogate dynamics, certified by a Lyapunov function, is preserved for the original dynamical system and vice versa. In the second part, we proposed a novel kernel-EDMD scheme for control-affine systems building upon arbitrarily-sampled control-state pairs and rigorously showed the first uniform (proportional) bounds on the full approximation error for control systems. We demonstrated that, as a consequence, stabilizing feedback laws designed for the data-driven surrogate models also ensure asymptotic stability w.r.t. the closed-loop of the original system building upon the proposed stability-analysis framework for kernel EDMD. We accompanied our findings by various numerical simulations.
Future work will be devoted to data-driven controller design leveraging the proposed highly-flexible sampling regime in combination with the novel proportional bounds derived in a non-restrictive setting. Herein, results from [18] and [11] may be leveraged for robustification and direct data-driven controller design. To alleviate the curse of dimensionality, future work considers leveraging low-dimensional structures by means of, e.g., tensor-train approximations [17], greedy methods for feature selection [48] or neural networks for dictionary learning [20].
@@ A.1. Error bounds
The first claim on the error between $\mathcal{K}$ and $\widetilde{\mathcal{K}}_{\lambda}$ follows by a combination of [28, Theorem 5.2] and Lemma 2.5. The proof of the second claim on $F-\widetilde{F}_{\lambda}$ is very similar to that of Corollary 2.8. ∎
@@ Appendix B The regularization operator
Similarly, one shows that $\langle f,R_{\mathcal{X}}^{\lambda}g\rangle=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}g_{\mathcal{X}}$. This shows $R_{\mathcal{X}}^{\lambda}=(R_{\mathcal{X}}^{\lambda})^{*}\,\geq\,0$.